Forecast.

Screen Shot 2023-11-19 at 12.09.17 PM.png

Details.

Telescope closed due to bad weather.

Requested AO for LuciDX.

Shark-Vis says No AO.

Screen Shot 2023-11-19 at 8.20.53 PM.png

08:00PM

Power on soul-luciDX.

ocam warm

09:07PM

Still closed.

Wait unitl 11:30PM

Still closed.

Observer called NO AO for tonight

-- JuanCarlosGuerra - 19 Nov 2023
Topic revision: r3 - 20 Nov 2023, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback